### v4

# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

Abstract
Knowledge of the effect of interventions, known as the treatment effect, is paramount for decision-making. Approaches to estimating this treatment effect using conditional average treatment effect (CATE) meta-learners often provide only a point estimate of this treatment effect, while additional uncertainty quantification is frequently desired to enhance decision-making confidence. To address this, we introduce two novel approaches: the conformal convolution T-learner (CCT-learner) and conformal Monte Carlo (CMC) meta-learners. The approaches leverage weighted conformal predictive systems (WCPS), Monte Carlo sampling, and CATE meta-learners to generate predictive distributions of individual treatment effect (ITE) that could enhance individualized decision-making. Although we show how assumptions about the noise distribution of the outcome influence the uncertainty predictions, our experiments demonstrate that the CCT- and CMC meta-learners achieve strong coverage while maintaining narrow interval widths. They also generate probabilistically calibrated predictive distributions, providing reliable ranges of ITEs across various synthetic and semi-synthetic datasets.
Code: https://github.com/predict-idlab/cct-cmc.
1 Introduction
"To act or not to act?" is a crucial question in decision-making across domains such as healthcare, policy formulation, and marketing [1, 2]. Today, decisions rely on the average results of studies such as clinical trials, A/B tests, or surveys. However, there is an increasing demand for more individualized interventions to optimize outcomes when a decision has to be made [1].
Specialized machine learning (ML) models offer solutions by estimating the conditional average treatment effect (CATE), predicting whether an intervention will be beneficial given the context [3, 4, 5, 6, 7]. However, CATE approaches typically provide single-point estimates, limiting their applicability in uncertain environments, especially in higher-risk settings such as healthcare [8, 9].
Uncertainty quantification of these CATE estimates and variations of individual treatment effect (ITE) addresses these limitations by capturing both the overall uncertainty in the estimate (epistemic) and the potential variation in ITEs (aleatoric). Traditionally, Bayesian approaches, such as Bayesian additive regression trees and Gaussian processes, have been used [10, 11] but they often suffer from under coverage and lack model-agnosticism, limiting their compatibility with modern ML models [12]. Alternatively, frequentist approaches, such as causal forests [3], only quantify the (epistemic) uncertainty around the CATE estimate, whereas a quantile regression approach to potential outcomes using the pinball loss function [13] only gives prediction intervals for the counterfactual outcomes and requires a Bonferroni type of correction, resulting in highly conservative intervals for ITE.
Conformal prediction (CP) [14] offers a promising alternative, providing distribution-free prediction intervals with finite-sample coverage guarantees under the exchangeability assumption. CP can be leveraged to generate prediction intervals with conservative coverage guarantees. Weighted CP accounts for the violation of the exchangeability assumption due to possible covariate shift during counterfactual inference [12]. Alternatively, one can alleviate the covariate shift by applying CP on the pseudo-outcomes of the two-step meta-learner [15]. While both approaches address calibration issues, they are often overly conservative and only generate prediction intervals, offering limited insights compared to predictive distributions which enable to construct decision-making systems by optimizing utility functions [16].
Contributions
This work introduces the conformal convolution T-learner (CCT-learner), a framework built on single-step meta-learners, that leverages convolutions and (weighted) conformal predictive systems (WCPS) [17, 18] to provide predictive distributions of ITE outcomes and less conservative prediction intervals. The CCT-learner outperforms existing prediction interval approaches for ITE by employing WCPS for counterfactual inference, addressing covariate shifts through propensity score-based weighting. The resulting calibrated distributions for treated and control outcomes are then convoluted to obtain the ITE distribution. Although our approach assumes conditionally independent predictive uncertainty of (potential) outcomes, we still have conservatively valid prediction intervals if the predictive uncertainty is positively correlated. Additionally, we argue that conditional independence of the (potential) outcomes aleatoric uncertainty is a needed assumption to provide calibrated predictive distributions and postulate this as the epsilon problem.
We demonstrate the CCT-learner’s effectiveness through simulations on three synthetic and three semi-synthetic datasets. As the convolution operation on nonparametric distributions during inference incurs significant computational cost, we additionally introduce the conformal Monte Carlo (CMC) T-learner that replaces the convolution with Monte Carlo sampling during training. This approach generates pseudo-outcomes for training a CPS, resulting in faster inference times. Simulations show that the CMC T-learner achieves coverage levels comparable to the CCT-learner with minimal efficiency loss. The introduced CMC framework allows adaptations to other meta-learners, such as the S-learner and X-learner [4]. We also establish the existence of a universal CCT- and CMC T-learner which converges to the true data-generating distribution as training data grows infinitely.
2 Background
Problem definition
In this work, we use the Neyman-Rubin [19] potential outcome framework to formulate and discuss our problem. We assume access to a sample of observations $\mathcal{D}=\{(Y_{i},X_{i},W_{i})\}^{n}_{i=1}$, with $(Y_{i},X_{i},W_{i})\overset{\mathrm{iid}}{\sim}\mathcal{P}$ with $\mathcal{P}$ the resulting joint distribution of the data generation process. Here, $Y_{i}\in\mathcal{Y}$ represents the continuous outcome of interest, $X_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ denotes the per-person feature values, and $W_{i}\in\{0,1\}$ is the binary treatment assignment. The joint distribution $\mathcal{P}$ is defined by following the covariate distribution and nuisance functions:
$\pi(x)$ $=\mathbb{P}(W=1|X=x)$ $\mu_{w}(x)$ $=\mathbb{E}(Y|X=x,W=w)$
$X$ $\sim\Lambda,$ $W$ $\sim\text{Bernoulli}(\pi(x))$
$Y^{w}$ $=\mu_{w}(X)+\varepsilon^{w},$ $w$$\in\{0,1\}$
$Y$ $=Y^{0}(1-W)+Y^{1}W$ $\tau(x)$ $=\mathbb{E}(Y^{1}-Y^{0}|X=x)$
$ITE$ $=Y^{1}-Y^{0}=\tau+\varepsilon_{a,ITE}$ $\varepsilon_{a,ITE}$ $=\varepsilon^{1}-\varepsilon^{0}$
where $\pi(x)$ denotes the propensity score that defines the probability of treatment assignment for an individual, $\mu_{w}(x)$ represents the expected (potential) outcome function of an individual for treatment assignment $w$, $\Lambda$ represents the marginal distribution of $X$, $Y^{w}$ represent the outcome when an individual receives treatment $w$, $\varepsilon^{w}$ are zero-mean random variables independent of $X$ and $W$, $Y$ is the observed outcome, $\tau(x)$ denotes the CATE function, $ITE$ the individual treatment effect, and $\varepsilon_{a,ITE}$ represent a zero-mean random variable. Next, we assume that the described data-generating process follows the following assumptions: (1) Consistency: $Y^{w_{i}}=Y_{i}$; (2) Unconfoundedness: $Y^{0},Y^{1}\perp W|X$ or equivalently $\varepsilon(0),\varepsilon(1)\perp W|X$; and (3) Positivity: $Pr[W=w|X=x]>0$. Under Assumption (1) and (2), $\mu_{w}$ becomes a potential outcome as $\mathbb{E}(Y^{w}|X)=\mathbb{E}(Y|X=x,W=w)$, aiding in estimating $\tau$ and defining the potential outcome $\mu_{w}$.
In this work, we aim to estimate ITE, defined as the difference between potential outcomes $Y^{0}_{i}$ and $Y^{1}_{i}$. However, as observing both outcomes for the same individual $i$ is impossible, often called the fundamental problem of causal inference, previous works focused on estimating the CATE function $\tau(x)$ [3, 4, 5, 6, 7]. When conditioning on the same covariate set, the optimal estimator for the CATE function also serves as the optimal estimator for the ITE function in terms of mean squared error [4]. However, as already stated, CATE approaches predominantly yield point estimates and often overlook quantifying uncertainty in predictions. While the intrinsic value of CATE estimates is undeniable, quantifying uncertainty surrounding these estimates is essential for robust decision-making in high-risk environments [8, 9]. Therefore, it is imperative to address both dimensions of predictive uncertainty, $\widehat{ITE}=\tau(X)+\varepsilon_{a,ITE}(X)+\varepsilon_{e,ITE}(X)$ with the epistemic uncertainty ($\varepsilon_{e,ITE}$) originating from model misspecification and finite samples, and the aleatoric uncertainty ($\varepsilon_{a,ITE}$) arising from the inherent variability of ITE among individuals with the same covariates. Incorporating a comprehensive understanding of these uncertainties is pivotal for advancing the reliability and applicability of heterogeneous treatment effect models.
CATE meta-learners
CATE estimation is frequently performed using CATE meta-learners. CATE meta-learners are frameworks where base learners are utilized to learn and perform CATE estimation [4]. Some examples are the T-learner, S-learner, X-learner [4], DR-learner [6], and R-learner [5]. We can categorize these learners into two groups: one-step plug-in (indirect) learners and two-step (direct) learners. One-step plug-in learners (S- and T-learners) are approaches that use two regression functions where the difference represents the CATE function and thus is estimated indirectly. In contrast, two-step learners (X-, DR-, and R-learners) output the CATE function directly. Three learners are used in this work: the T-, S-, and X-learner. In the binary treatment case, the T-learner fits two different base learners $\mu_{0}$ and $\mu_{1}$ on data without treatment and with treatment, respectively. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}_{1}(X)-\hat{\mu}_{0}(X)$. In contrast, the S-learner fits a single base learner $\mu$ on $[X,W]$ with $W$ the treatment variable. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}(X,W=1)-\hat{\mu}(X,W=0)$. The X-learner starts as a T-learner by first fitting $\mu_{0}$ and $\mu_{1}$, however, afterwards a direct estimate of $\tau$ is fitted with $X$ as input and $W(Y-\hat{\mu}_{0}(X))+(1-W)(\hat{\mu}_{1}(X)-Y)$ as target.
Conformal prediction
Conformal prediction (CP) is a model-agnostic framework that gives an implicit confidence estimate in a prediction by generating prediction sets at a specified significance level $\alpha$ [14]. Notably, this framework allows for (conservatively) valid non-asymptotic confidence predictors under the exchangeability assumption, assuming training/calibration data exchangeability with test data, a slightly less stringent assumption than data independence. The prediction sets in CP are formed by comparing nonconformity scores of examples that quantify how unusual a predicted label is, i.e., these scores measure the disagreement between the prediction and the actual target.
More specifically, we define a prediction interval $\hat{C}(x_{n+1})$, for test object $x_{n+1}\in\mathbf{X}$, by computing following conformity scores $R_{i}^{y}$, based on conformity measure $A$, for each $y\in\mathbb{R}$, using training data $z_{i}=(x_{i},y_{i}),i=1,...,n$:
$R_{i}^{y}=A(z_{i},z_{1:n\setminus i}\cup(x_{n+1},y)),\quad i=1,...,n$ (1)
and
$R_{n+1}^{y}=A((x_{n+1},y),z_{1:n}).$ (2)
The label $y$ is included in prediction interval $\hat{C}(x_{n+1})$ if,
$\frac{|i=1,...,n+1:R_{i}^{y}\geq R_{n+1}^{y}|}{n+1}>\alpha.$ (3)
This procedure, termed full or transductive conformal prediction, is computationally intensive. Therefore, a more applicable variant of full CP, called inductive or split CP (ICP) [20], offers computational efficiency and enables the integration of CP with ML algorithms, such as neural networks and tree-based models. In ICP, the training sequence $(x,y)=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ is portioned into a proper training sequence $\{(x_{1},y_{1}),...,(x_{m},y_{m})\}$ and a calibration sequence $\{(x_{m+1},y_{m+1}),...,(x_{n},y_{n})\}$. We first train a regression model $\hat{y}(x)$ using the proper training sequence. Then, we generate nonconformity scores $R_{i}$ for data points $(x_{i},y_{i})$ where $m<i\leq n$. These scores, which can be calculated using metrics like absolute error $R_{i}=|y_{i}-\hat{y}_{i}|$, are subsequently sorted in descending order, resulting in a ranking: $R^{*}_{1},...,R^{*}_{n-m}$. For a new test object $x_{n+1}$ and point prediction $\hat{y}_{n+1}$, ICP outputs then a prediction interval for a desired target coverage of $1-\alpha$:
$[\hat{y}_{n+1}-R^{*}_{s},\hat{y}_{n+1}+R^{*}_{s}]$ (4)
where $s=\lfloor\alpha(n-m+1)\rfloor$.
Although the conventional ICP yields calibrated prediction intervals, it only offers marginal coverage and constant interval sizes for different covariates. As we want smaller intervals for confident predictions and larger intervals for more "difficult" examples, Mondrian CP [21] and conformalized quantile regression (CQR) [22] extensions of the ICP procedure are proposed. Another adaptation of ICP for establishing more adaptive prediction intervals involves employing normalized nonconformity measures [23]. These normalized measures typically use the product of the absolute error $|y-\hat{f}(x)|$ and the reciprocal of a 1D uncertainty estimate $\frac{1}{\hat{u}(x)}$, expressed as $R=\frac{|y-\hat{f}(x)|}{\hat{u}(x)}$, where $\hat{u}(x)$ may be generated by a distinct model which estimates $|y-\hat{f}(x)|$. Alternatively, $\hat{u}(x)$ can also be obtained using, e.g., the estimated variance from Monte Carlo dropout or an ensemble model [24]. Following a similar procedure as before, we estimate $\hat{u}(x)$ and adjust the predicted output accordingly, resulting in a more adaptive prediction interval:
$[\hat{y}_{n+1}-\hat{u}(x)R^{*}_{s},\hat{y}_{n+1}+\hat{u}(x)R^{*}_{s}]$ (5)
For a more comprehensive exploration of CP, its extensions and applications, we refer the reader to Vovk et al. [14], Angelopoulos and Bates [24], Fontana et al. [25], Manokhin [26].
Conformal predictive systems
Recently, Vovk et al. [17] proposed conformal predictive systems (CPS), an extension of full CP, that allow deriving predictive distributions and enjoys provable properties of validity, in the sense of probabilistic calibration, under IID assumption [17]. CPS produces conformal predictive distributions (CPD) as p-values arranged into a probability distribution function [17], created with the help of conformity measures. Vovk et al. [17] define a CPS as a function that is both a smooth conformal transducer and a randomized predictive system (RPS). The conformal transducer is another way to represent a conformal predictor, i.e., a confidence transducer that can be seen as a function that maps each input sequence into a sequence of p-values [14]. Formally, we define the full smoothed conformal transducer defined by conformity measure $A$ as,
$Q(z_{1},...,z_{n},(x,y),\phi):=\sum_{i=1}^{n+1}[R_{i}^{y}<R_{n+1}^{y}]\frac{1}{n+1}+\sum_{i=1}^{n+1}[R_{i}^{y}=R_{n+1}^{y}]\frac{\phi}{n+1}$
where $(z_{1},...,z_{n})$ is the training sequence, $\phi\in[0,1]$, $x_{n+1}$ is a test object, and for each label $y$ the corresponding conformity score $R_{i}^{y}$ is defined as
$\begin{split}R_{i}^{y}&:=A(z_{1},...,z_{i-1},z_{i+1},...,z_{n},(x_{n+1},y),z_{i}),\qquad i=1,...,n\\ R_{n+1}^{y}&:=A(z_{1},...,z_{n},(x_{n+1},y)).\end{split}$
The RPS is then a function $Q:\textbf{Z}^{n+1}\times[0,1]\rightarrow[0,1]$ that satisfies three requirements: (1) the function $Q(z_{1},...,z_{n},(x,y),\phi)$ needs to be monotonically increasing both in $y\in\mathbb{R}$ and $\phi\in[0,1]$; (2) $\lim_{y\rightarrow-\infty}Q(z_{1},...,z_{n},(x,y),0)=0$ and $\lim_{y\rightarrow\infty}Q(z_{1},...,z_{n},(x,y),1)=1$; and (3) it must be probabilistically calibrated, $\forall\alpha\in[0,1]:\mathbb{P}\{Q(z_{1},...,z_{n},z_{n+1},\tau)\leq\alpha\}=\alpha.$
A smoothed conformal transducer is an RPS if the IID assumption is satisfied and if it is defined by a monotonic conformity measure [17].
CPS has been adapted to become more computationally efficient, i.e., split conformal predictive systems (SCPS) [27], by building upon ICP. Here, the p-values are created by a split conformity measure that is an element of a family of measurable functions $A_{m}:\{(X_{1},y_{1}),...,(X_{m},y_{m},(X,y)\}\rightarrow\mathbb{R}\cup\{-\infty,\infty\}$. This measure needs to be isotonic and balanced. Examples of such a measure are the residual error $y-\hat{f}(x)$ and normalized residual error $\frac{y-\hat{f}(x)}{\hat{u}(x)}$. The SCPS is a function that is both a split-smoothed conformal transducer and an RPS. The split conformal transducer determined by a split conformity measure $A$ is defined as
$Q(z_{1},...,z_{n},(x,y),\tau):=$ $\sum_{i=m+1}^{n}[R_{i}<R^{y}]\frac{1}{n-m+1}+\sum_{i=m+1}^{n}[R_{i}=R^{y}]\frac{\tau}{n-m+1}$
$+\frac{\tau}{n-m+1}$
where conformity scores $R_{i}$ and $R^{y}$ are defined by
$\begin{split}R_{i}&:=A(z_{1},...,z_{m},(x_{i},y_{i})),\qquad i=m+1,...,n,\\ R^{y}&:=A(z_{1},...,z_{m},(x,y)),\qquad y\in\mathbb{R}.\end{split}$
Vovk et al. [27] proved that the split smoothed conformal transducer is an RPS if the IID assumption is satisfied and the conformity measure is isotonic and balanced.
Conformal prediction and conformal predictive systems under covariate shift
As mentioned, CP and CPS validity relies on the exchangeability and IID assumptions, respectively. However, in counterfactual inference, except in randomized control trials (RCTs), these assumptions are often violated due to covariate shift. Treatment assignment probabilities, depending on covariates, cause this shift, affecting the modeling of counterfactual outcomes.
Fortunately, Tibshirani et al. [28] extended CP beyond the exchangeability assumption, proposing weighted CP (WCP), a weighted version of CP to deal with covariate shifts where the likelihood ratio between the training $P_{X}$ and test $\tilde{P}_{X}$ is known. WCP reweights each nonconformity score by a probability proportional to the likelihood ratio $\frac{d\tilde{P}_{X}(x_{i})}{dP_{X}(x_{i})}$. Lei and Candès [12] implemented it for counterfactual inference, next to extending it to CQR and giving an upper bound to the conservative validity guarantee of WCP. For CPS, Jonkers et al. [18] proposed weighted CPS (WCPS), similar to WCP, to allow for probabilistically calibrated predictive distributions under covariate shift.
3 Conformal Convolution and Conformal Monte Carlo Meta-Learner
The main concept behind both conformal convolution (CC) and conformal Monte Carlo (CMC) T-learner is to start from counterfactual inference. They aim to generate probabilistically calibrated predictive distributions to infer the ITE predictive distribution. The CCT- and CMC T-learner methods are illustrated in Figure 1, with descriptions provided in Algorithm 1 and 2.
Figure 1: Illustration of CCT- and CMC T-learner algorithms.
Algorithm 1 Fit and inference of CCT-learner
1:  Assume: $\phi\sim Uniform(0,1)$, $y\in\mathbb{R}$
2:  Input: training covariates $X_{tr}$, labels $y_{tr}$, treatment indicator $W_{tr}$, propensity score function $\pi(x)$, and test object $x_{test}$
3:  $X_{tr,0}$, $y_{tr,0}$:= $X_{tr}[W_{tr}=0]$, $y_{tr}[W_{tr}=0]$
4:  $X_{tr,1}$, $y_{tr,1}$:= $X_{tr}[W_{tr}=1]$, $y_{tr}[W_{tr}=1]$
5:  $\textbf{Q}_{Y^{0}}(x_{test},y,\phi):=$$\textbf{WCPS}(X_{tr,0}$, $y_{tr,0}$, $x_{test}$, $\frac{1}{1-\pi(x)}$,$y$,$\phi)$
6:  $\textbf{Q}_{Y^{1}}(x_{test},y,\phi):=$$\textbf{WCPS}(X_{tr,1}$, $y_{tr,1}$, $x_{test}$, $\frac{1}{\pi(x)}$, $y$, $\phi)$
7:  $\textbf{Q}_{ITE}(x_{test},y,\phi)=$$\textbf{Q}_{Y^{1}}*\textbf{Q}_{-Y^{0}}(x_{test},y,\phi)$see Eq. 6}
Both methods deal with covariate shifts using WCPS. These shifts occur because the control and treatment models are trained and calibrated solely on observed outcomes, leading to different covariate distributions between treated and control groups due to confounding variables. To achieve calibrated predictive distributions for outcomes under control $Y^{0}$ and treatment $Y^{1}$, we weight each conformity scores proportional to $\frac{1}{1-\pi(x)}$ and $\frac{1}{\pi(x)}$ since
$w_{0}(x)$ $=\frac{dP_{X}(x)}{dP_{X|W=0}(x)}=\frac{dP_{X}(x)\mathbb{P}(W=0)}{dP_{X}(x)(1-\pi(x))}\propto\frac{1}{1-\pi(x)}$
$w_{1}(x)$ $=\frac{dP_{X}(x)}{dP_{X|W=1}(x)}=\frac{dP_{X}(x)\mathbb{P}(W=1)}{dP_{X}(x)\pi(x)}\propto\frac{1}{\pi(x)}$
where $P_{X}(x)$ denotes the covariate distribution of the entire population, $P_{X|W=1}(x)$ and $P_{X|W=0}(x)$ represent the covariate distribution of the treated and control population, and $\mathbb{P}(W=1)$ and $\mathbb{P}(W=0)$ are the marginal probabilities of receiving treatment or not, respectively. These weights align with those used by Lei and Candès [12] for counterfactual inference with WCP and are similar to the inverse probability weighted (IPW) estimation of average causal effects [29].
From the WCPS, we infer the potential outcomes predictive distributions, $Q_{Y^{0}}$ and $Q_{Y^{1}}$, which are probabilistically calibrated. Our interest lies in the ITE, $Y^{1}-Y^{0}$, so we perform a convolution operation on the treatment and the negated control predictive distributions, defined as:
[…]
This integral is intractable and requires numerical integration techniques, making the WCPS computationally expensive on top of the already expensive full WCPS. Therefore, we use a split WCPS approach in practice, defining the convolution as follows:
$\begin{split}Q_{ITE}(x_{test},\pi,y,\phi)&=\sum_{i=1}^{N_{cal,Y^{1}}}\sum_{j=1}^{N_{cal,Y^{0}}}[\{q_{i,Y^{1}}(x_{test})-q_{j,Y^{0}}(x_{test})\}<y]p_{i}^{w_{1}}(x_{test})p_{j}^{w_{0}}(x_{test})\\ &+[\{q_{i,Y^{1}}(x_{test})-q_{j,Y^{0}}(x_{test})\}=y]p_{i}^{w_{1}}(x_{test})p_{j}^{w_{0}}(x_{test})\phi\\ &+[1-\{1-p_{test}^{w_{1}}(x_{test})\}\{1-p_{test}^{w_{0}}(x_{test})\}]\phi\end{split}$ (6)
where
$p_{i}^{w}(x_{test})=\frac{w(x_{i})}{\sum_{j=1}^{n}w(x_{j})+w(x_{test})},\qquad p_{test}^{w}(x_{test})=\frac{w(x_{test})}{\sum_{j=1}^{n}w(x_{j})+w(x_{test})},$ (7)
$N_{cal,Y^{1}}$ and $N_{cal,Y^{0}}$ are the calibration set sizes for the split WCPS of outcomes under treatment and control, $q_{i,Y^{1}}(x_{test})$ and $q_{j,Y^{0}}(x_{test})$ are the $i$th and $j$th right-continuous points of $Q_{Y^{1}}(x_{test},y,\phi=0)$ and $Q_{Y^{0}}(x_{test},y,\phi=0)$, and $\pi$ represents the propensity score function.
While this split CCT-learner is way more computationally efficient, the convolution applied during inference, see Figure 1, is quite impractical in certain cases. Therefore, we propose the CMC T-learner, which replaces the convolution performed during inference with Monte Carlo sampling during training. The CMC framework samples from the fitted (potential) outcome distributions, creating Monte Carlo (MC) samples of the (potential) outcomes. Subsequently, these MC samples are utilized to artificially generate ITE samples. To generate these ITE samples, each example in the calibration set is sampled a specified number of times, producing multiple MC ITE samples for the same covariates. These ITE samples, treated as ground truth, are then employed alongside a meta-learner functioning as a CATE estimator to define a CPS, ultimately yielding conditional ITE predictive distributions.
We propose two methods for generating MC ITE samples: (1) The MC ITE method independently samples from both conditional potential outcome predictive distributions, subtracting the potential outcome under control $\hat{Y}^{0}$ from the one under treatment $\hat{Y}^{1}$; (2) The pseudo-MC (pMC) ITE method adapts this by conditioning on the observed treatment value, calculating the ITE sample as the difference between the observed outcome $Y$ and the conditional potential outcome predictive distribution under control $\hat{Y}^{0}$, and vice versa when treatment is not administered. Since the two approaches sample from different covariate distributions, they require a different weighting to be applied to WCPS. While the MC requires the same weighting scheme as in the CCT-learner, the pMC procedure requires a weighting of $\frac{\pi(x)}{1-\pi(x)}$ and $\frac{1-\pi(x)}{\pi(x)}$ for the counterfactual outcome of the treated and untreated subjects. Algorithm 3 provides a comprehensive overview of this sampling scheme.
We also introduce two variants of CMC meta-learners: CMC S- and X-learners, that follow the same process for generating conditional ITE predictive distributions, differing slightly in modeling nuisance predictive distributions and leveraging MC ITE samples. Details are provided in Appendix B.
Both CCT- and CMC meta-learners return conditional predictive distributions of the ITE. Prediction intervals of confidence level $1-\alpha$ are formed by taking the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles from the predictive distribution as the lower and upper bounds, respectively. Our method produces symmetric, two-sided intervals with equally distributed probability density. In contrast, the approaches by Lei and Candès [12] and Alaa et al. [15] generate asymmetric intervals, offering an efficiency advantage due to smaller interval lengths [22]. However, these asymmetric intervals can be misleading in practical applications as they might intuitively imply equal probability density on both sides of the interval.
3.1 The Epsilon Problem
In addition to introducing the CCT- and CMC meta-learners, we propose to add a fourth assumption to the three assumptions (see Section 2) on the underlying data-generating processes to perform CATE estimation to define its relation with the validity of predictive inference on ITEs. This fourth assumption covers the relation between noise distributions under treatment $\varepsilon(1)$ and control $\varepsilon(0)$. If we want to use the CP framework for performing valid predictive inference, we need to know their joint distribution. As we never jointly observe $\varepsilon(0)$ and $\varepsilon(1)$, and can thus not learn their joint distribution, we need to assume how these two noise distributions are related. The literature related to CATE estimation often assumes, when creating data-generating processes, that these noise terms are zero-mean random variables and independent of X and W  [4, 5, 6, 7]. Some works [5, 6, 7] assume that the treatment effect is deterministic and thus, in consequence, that the two noise terms are identical for the same example, whereas Künzel et al. [4] assumes that the noise distributions are fully independent in their synthetic simulation experiments. The limited literature concerning predictive inference of ITE  [10, 11, 12, 15] mostly assume that the noise distributions are fully independent and set up their experiments in the same way. In the following paragraphs, we will therefore discuss three possible characteristics of the noise terms’ joint distribution and their impact on the validity of predictive intervals and distributions of ITE. While the true nature of these joint noise distributions is, in part, a philosophical question, it requires further attention from the research community and domain experts.
$\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$: In instances where the two noise distributions are mutually independent, we anticipate reasonable validity of predictive distributions with the CCT- and CMC T-learner, and, at the very least, the capability to generate conservative prediction intervals for the other CMC meta-learners.
$\mathbf{\boldsymbol{\varepsilon(0)=\varepsilon(1)}}$: In the scenario where the two noise terms are identical for each sample, a common assumption in most CATE simulation experiments, it is essential to acknowledge that the noise distributions become nullified when inferring ITE. This way, the ITE becomes entirely deterministic. Here both the CCT- and CMC meta-learners over account for uncertainty since the noise distribution (i.e., the aleatoric part) is shared in both (potential) outcomes. To address this, we suggest modeling the epistemic uncertainty distribution of two potential outcome models as future work. Assuming these epistemic uncertainty distributions are independent of each other, the convolution and the MC sampling become more valid. Nevertheless, it is important to note that here the CCT- and CMC meta-learners will continue to provide conservatively valid prediction intervals.
[…]
In this paper, simulation experiments, presented in Figure 3, further support these statements. Note that these assumptions also affect other predictive inference approaches for the ITE and are, therefore, not limited to CCT- and CMC alone.
3.2 Universal Consistent CCT- and CMC T-Learner
The introduction of the fourth assumption allows us to prove the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data-generating distribution as $n\rightarrow\infty$ under any date-generating distribution, which we formalize in Theorem 1.
Theorem 1.
Assuming independent (potential) outcome noise distributions $\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$, and that the object space $\mathcal{X}$ is standard Borel, a universal consistent CCT- and CMC T-learner exists (see Algorithm 1 and 2).
The proof for Theorem 1 is given in Appendix A. In short, it uses Theorem 3 of [30], stating that a universal consistent CPS exists and that the convolution of two independent weakly converging sequences weakly convergence to the independent joint distribution.
4 Experiments
4.1 Experimental Setup
Similar to Nie and Wager [5] and Alaa et al. [15], we evaluated our CCT- and CMC T-learners through six different experiments: three using synthetic data (see Section D.1) and three using semi-synthetic data (see Section D.2). The experiments assessed coverage and efficiency of prediction intervals of the CCT-learner and (pseudo) CMC T-learner against WCP (exact, inexact, and naive) [12], CM-Learner (DR) [15], and a CPS Oracle fitted as a standard regression problem on the ITE (see Appendix C). Additionally, we evaluate the probabilistic calibration on six different synthetic dataset settings. All learners utilized the scikit-learn RandomForestRegressor with default hyperparameters. The CMC learners always employed 100 MC samples. Although experiments with other base learners like gradient-boosting machines were conducted, they were omitted as their results were similar to those of random forests. Additional results of the CMC S- and X-learners are included in Appendix E.
4.2 Results
(a) Probability Integral Transform (PIT)
(b) 90% ITE prediction interval.
Figure 2: Simulation results for setups A and B from Alaa et al. [15].
Figure 3: Simulation results for setups A, B, C, and D from Nie and Wager [5], with increasing correlation between the noise terms of potential outcome under treatment $Y^{1}$ and control $Y^{0}$. The vertical lines represent the variation in the different simulation runs and cover an interval with the results of 95% of the runs.
Figure 4: Results on 77 settings of the ACIC2016 dataset for the 90% ITE prediction interval.
Coverage and efficiency of prediction intervals
On synthetic datasets, the CCT- and CMC meta-learners consistently demonstrated slightly conservative coverage across various simulations while maintaining high efficiency, indicated by small average interval widths. In contrast, the inexact WCP, while more efficient in some cases, often did so at the cost of significant under coverage. The CM DR learner was consistently more conservative than our proposed methods. Other WCP approaches, with proven conservative validity, were extremely conservative across all confidence levels and experiments, likely due to their over-conservative design, employing either Bonferroni correction or double conformal inference procedure [12].
Comparing the CCT- and (p)CMC T-learners, both showed similar coverage levels across all experiments. The pCMC T-learner produced less conservative intervals than the other two approaches; however, this did not necessarily translate to greater efficiency compared to the CMC T-learner. The CCT-learner consistently produced more efficient intervals across all datasets, possibly because it can use more samples for the potential outcome models as the (p)CMC learners require an additional data split. Applying a cross-fitting procedure might mitigate this effect for the (p)CMC learners, a possibility we leave for future work. Additionally, the pCMC showed slightly greater computational efficiency, needing one less variable per MC sample than the CMC learner.
Within the CMC meta-learners, performance differences were small and dependent on the data-generating process. For example, in setup D from Nie and Wager [5], where treatment and control groups have unrelated outcomes, the T-learner was slightly more efficient than the CMC S- and X-learners, possibly due to the absence of synergies for jointly learning both outcomes.
For the three semi-synthetic datasets, we performed experiments with covariates from real-world data. As here no propensity scores are provided, these resemble a realistic scenario. We estimate the propensity scores using logistic regression. The most extensive experiment was conducted on the 77 settings of the ACIC2016 dataset [31], which contain 58 covariates (3 categorical, 5 binary, 27 counts, 23 continuous) with varying complexity of outcome surfaces, degree of confounding, overlap, and treatment effect heterogeneity. Results, depicted in Figure 4, show that our proposed approaches provide conservatively valid prediction intervals across all settings, with significant conservative coverage in most settings, though less than the CM DR-learner and valid WCP approaches, likely due to violations in the assumption of conditional independence of the noise distributions. Regarding efficiency, the CCT-learner performed best for the reasons previously mentioned. As can be seen in Figure 11, a similar conclusion can be drawn for the other semi-synthetic experiments.
Performance under different relations between noise distributions
The relationship between the noise distributions of the two potential outcomes significantly affects coverage and the validity of interval predictions and predictive distributions, as evidenced in our experiments where we incrementally increased correlations between the two noise terms. Figure 3 shows that all methods exhibit more overconfidence as correlation increases starting from zero. When correlation becomes negative, both the CCT- and CMC-learners lose some validity, though not as much as the inexact WCP. The CM DR-learner and valid WCP approaches, while experiencing decreased coverage, remain valid. While this is one of the limitations of CCT- and CMC-learners, as this is an untestable condition, one could argue if such a scenario happens in reality.
Probabilistic calibration of CCT- and CMC T-learner
We assessed the probabilistic calibration of the predictive distribution by plotting the probability integral transform (PIT) values across different experiments and simulation runs. A probabilistic forecaster is considered calibrated if it produces uniformly distributed PIT values on the test set [31]. Figure 2(a) shows these PIT values for the synthetic experiments from Alaa et al. [15], indicating slight conservative coverage for both the CCT-learner and CMC T-learner, though minimal.
The learners are almost perfectly calibrated for the synthetic datasets from Nie and Wager [5], as shown in Figure 9. In general, the slight conservative coverage could be due to dependent epistemic uncertainty from shared components between outcomes under treatment and control. Setup D was perfectly calibrated with unrelated outcomes under treatment and control (see Figure 9), supporting this hypothesis. Additionally, Figure 10 illustrates the effect of replacing WCPS with CPS, resulting in uncalibrated learners, except for Setup B, which represents a randomized control trial.
5 Conclusion
In this work, we present the CCT- and CMC meta-learners for estimating the predictive distribution of CATE estimation to reveal information about the underlying ITE, enhancing decision-making confidence. To the best of our knowledge, our frameworks are the first frequentist approaches to offer an ITE predictive distribution. They outperform the state-of-the-art by demonstrating strong experimental coverage, maintaining narrow interval widths and achieving probabilistic calibration on six different datasets, while enabling the generation of predictive non-parametric distributions. We also highlighted how different assumptions about the noise distribution of potential outcomes can significantly impact the coverage of ITE predictive inference.
Acknowledgement
Part of this research was supported through the Flemish Government (AI Research Program). Jarne Verhaeghe is funded by the Research Foundation Flanders (FWO, Ref. 1S59522N).
Appendix A Proof: Existence Universal Consistent CCT- and CMC T-Learner
The following section presents and proves Theorem 1, which states the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data generating distribution as $n\rightarrow\infty$ under any data-generating distribution. We introduce here Definition 1, taken in verbatim from [30], which lays out our notion of consistency for CPS.
Definition 1 (in verbatim Vovk [30]).
A conformal predictive system $Q$ is consistent for a probability measure $\mathcal{P}$ on Z := X × R if, for any bounded continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$,
[…]
in probability, where:
  • $Q_{n}$ is the predictive distribution $Q_{n}:y\mapsto Q(z_{1},...,z_{n},x_{test},\phi)$ output by Q as its forecast for the label $y_{test}$ corresponding to the test object $x_{test}$ based on the training data $(z_{1},...,z_{n})$, where $z_{i}=(x_{i},y_{i})$ for all $i$.
  • $\mathbb{E}_{\mathcal{P}}(f|X_{test})$ is the conditional expectation of $f(y)$ given $x=x_{test}$ under $(x,y)\sim\mathcal{P}$.
  • $z_{i}=(x_{i},y_{i})\sim\mathcal{P},i=1,...,n$, $z_{test}\sim\mathcal{P}$, and $\phi\sim Uniform(0,1)$, are assumed all independent.
Theorem 1.
A universal consistent CCT- and CMC T-learner exists (see Algorithm 1 and 2). Assuming independent (potential) outcome noise distributions: $\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$, and the object space $\mathcal{X}$ is standard Borel.
Proof.
The philosophy behind the MC simulations in the CMC T-learner is to not perform a convolution of both outcome predictive distributions, as this would quickly become an expensive operation that would need to be performed for every example during inference. However, to prove that there exists a universal consistent, we replace this MC sampling in Algorithm 2 with the convolution operation,
[…]
and thus becoming a CCT-learner, which is equivalent to the CMC T-learner approach asymptotically.
As we consider the existence of a universal consistent learner, we no longer need to worry about the covariate shift. We can replace the WCPS with CPS in both the CCT- and CMC T-learner. This is equivalent to setting the propensity score to $0.5$ for every covariate.
By Theorem 3 in [30], we know that there exists universal CPS if the object space $\mathcal{X}$ is standard Borel, meaning that for any probability measure $\mathcal{P}$ on $Z$, the CPS is consistent. Consequently, there exists a universal CPS that can model the true data-generating distribution outcome under treatment $Y^{1}$ and control $Y^{0}$. These distributions are weakly converging distributions. Since we assume that the predictive distributions of the outcome are independent,
$Y^{0}\perp\!\!\!\perp Y^{1}|X,$ (10)
the outcome variables are independent conditionally on the covariate set $X$. Then, by the definition of convolution and the fact that $ITE=Y^{1}-Y^{0}$, the convolution of the true conditional cumulative distribution of outcome under treatment $F_{Y^{1}|X}$ and outcome under negative control $F_{-Y^{0}|X}$ equals the true conditional cumulative ITE distribution,
[…]
Since the convolution of two independent weakly converging sequences converges weakly to the independent joint distribution (Theorem 105, McLeish [32]), a universal consistent CCT-learner exists and consequently also a CMC T-learner. ∎
Appendix B Conformal Monte Carlo Meta-Learners
B.1 T-learner
Algorithm 2 Fit and inference of CMC T-learner
1:  Assume: $\phi\sim Uniform(0,1)$, $y\in\mathbb{R}$
2:  Input: training covariates $X_{tr}$, labels $y_{tr}$, treatment indicator $W_{tr}$, propensity score function $\pi(x)$, MC samples $n$, and test object $x_{test}$
3:  $X_{tn},X_{cal},y_{tn},y_{cal},W_{tn},W_{cal}=$split$(X_{tr},y_{tr},W_{tr})$
4:  $X_{..,0}$, $y_{..,0}$:= $X_{..}[W_{..}=0]$, $y_{..}[W_{tr}=0]$
5:  $X_{..,1}$, $y_{..,1}$:= $X_{..}[W_{..}=1]$, $y_{..}[W_{tr}=1]$
6:  $\textbf{Q}_{Y^{0}}(x,y,\phi):=$$\textbf{WCPS}(X_{tn,0}$, $y_{tn,0}$, $x$, $\frac{1}{1-\pi(x)}$,$y$,$\phi)$
7:  $\textbf{Q}_{Y^{1}}(x,y,\phi):=$$\textbf{WCPS}(X_{tn,1}$, $y_{tn,1}$, $x$, $\frac{1}{\pi(x)}$, $y$, $\phi)$
[…]
9:  $\textbf{Q}_{ITE}(x_{test},y,\phi)=\textbf{CPS}(repeat(X_{cal}$, $n$  times)))), $\textbf{ITE}_{MC},x_{test},y,\phi)$
Algorithm 3 (Pseudo) Monte Carlo ITE sampling
1:  Input: data $X$, target $y$, treatment $W$, predictive distributions $Q_{Y^{0}}$ and $Q_{Y^{1}}$, MC samples $n$
2:  $ITE_{MC}=zeros(len(X)\cdot n)$
3:  for $i=1$ to $len(X)$ do
4:     if Pseudo CMC then
5:        if $W_{i}==0$ then
6:           Sample $u_{1}\sim uniform(0,1,n$ times)
7:           $ITE_{MC,i\cdot n:(i+1)n}$ = $\textbf{Q}^{-1}_{Y^{1}}(X[i],u_{1})-y[i]$
8:        else
9:           Sample $u_{0}\sim uniform(0,1,n$ times)
10:           $ITE_{MC,i\cdot n:(i+1)n}$ = $y[i]$ $\textbf{Q}^{-1}_{Y^{0}}(X[i]$, $u_{0})$
11:        end if
12:     else
13:        Sample $u_{0}\sim uniform(0,1,n$ times)
14:        Sample $u_{1}\sim uniform(0,1,n$ times)
15:        $ITE_{MC,i\cdot n:(i+1)*n}$ = $\textbf{Q}^{-1}_{Y^{1}}(X[i],u_{1})-\textbf{Q}^{-1}_{Y^{0}}(X[i],u_{0})$
16:     end if
17:  end for
18:  return $ITE_{MC}$
B.2 S-Learner
The CMC S-learner closely resembles the T-learner algorithm with a notable distinction: it employs an S-learner as a CATE estimator and a single regression model for modeling the nuisance parameters. Consequently, we can only use split WCPS (SWCPS). This approach offers the advantage of leveraging a single model, thereby increasing the availability of samples to learn common components of the outcome. However, a drawback with this method arises from shared epistemic uncertainty, accounted for twice, potentially compromising the validity of the ITE predictive distributions. Nonetheless, the method remains capable of furnishing conservatively valid prediction intervals even in this scenario, as we observe in the simulation experiments.
B.3 X-Learner
The procedure for the CMC X-learner mirrors that of the CMC-T-learner, with a key distinction: instead of employing the T-learner as the CATE estimator, we incorporate an extra estimator by fitting a regression model on the generated MC ITE samples. A notable advantage is that the ITE regression model, serving as the CATE estimator, can utilize a reduced set of covariates compared to the input data for the nuisance models. This becomes particularly advantageous in scenarios where only a limited covariate set is accessible during inference. However, a potential drawback is the likelihood of sacrificing some level of validity, as the MC ITE samples derived from the nuisance models may introduce epistemic variability. Consequently, this variability might contribute to widening the predictive distribution. Nevertheless, despite this potential trade-off, the approach still enables the generation of conservatively valid prediction intervals, as we observe in the simulation experiments later presented in this work.
Appendix C Literature Overview
This section provides a more comprehensive overview of the literature on predictive inference ITE and quantifying related predictive uncertainty. Uncertainty quantification is a crucial component for treatment effect estimations for reliable decision-making in high-risk environments, such as healthcare [8]. In the past, most proposals for uncertainty quantification in ITE estimation used Bayesian approaches, such as Bayesian additive regression trees (BART) [10] and Gaussian processes [11]. The problem with these Bayesian approaches is that they are often uncalibrated without knowing the underlying data-generating process. These Bayesian approaches are also model-specific and are not easily transferred to modern ML approaches.
Recently, some works have tried to resolve this issue by using the CP framework  [12, 15], which can provide non-asymptotic distribution-free coverage guarantees and thus resolves the significant shortcomings of these Bayesian approaches. Lei and Candès [12] was the first work to propose a conformal inference approach to produce reliable prediction intervals for ITEs. More specifically, they proposed using WCP to account for covariate shifts as the distributions of covariates for treated and untreated subjects can differ from the target population. They proposed three variants of WCP ITE algorithms: (1) Naive WCP, which combines two potential outcome interval estimates, estimated by applying both WCP and a Bonferroni correction to get an ITE prediction interval; (2) Exact nested WCP, which first generates plug-in estimates of ITEs by using WCP and uses these estimates to apply a regular CP; and (3) Inexact nested WCP which replaces the second CP step in Exact WCP with quantile regression which does not provide any coverage guarantees. A significant drawback of the exact and inexact WCP approaches is that they require refitting the model, at least partially, for different confidence levels for the prediction intervals. Additionally, all the proposed approaches based on WCP only provide intervals as uncertainty quantification.
Alaa et al. [15] propose conformal meta (CM)-learners, which is a model-agnostic framework for issuing predictive intervals for ITEs that applies the inductive CP on top of CATE meta-learners. These CATE meta-learners are model-agnostic approaches that estimate the CATE by combining nuisance estimates and incorporating some inductive biases. These CM-learners are applied to a particular class of meta-learners, direct learners, which use a two-step procedure where nuisance estimators are fitted on the outcome in the first step. In the second step, these nuisance estimators are used to generate a pseudo-outcome in which the covariates are regressed on to get point estimates of the CATE; examples of such learners are X-, DR-, and IPW-learners. To construct prediction intervals of the ITE, Alaa et al. [15] propose to use these pseudo outcomes to generate conformity scores together with the estimates of the meta-learner, i.e., CATE estimator. However, their proposed framework requires more stringent assumptions to be valid than the classical ones for estimating CATE. They developed the so-called stochastic ordering framework and showed that CM-learners are marginally valid if their pseudo-outcome conformity scores are stochastically larger than the "oracle" conformity scores evaluated on the unobserved ITEs. However, it is difficult to translate this into practice, as the only way to verify this assumption is to observe the true ITEs, which is impossible. Additionally, their method is limited to only producing prediction intervals of ITE.
Appendix D Experimental setups
D.1 Synthetic Experiments
D.1.1 Alaa Synthetic Data Experiments
The first synthetic data experiments follows the setup proposed by Alaa et al. [15]. Two setups are used: A and B. Each setup uses 5000 samples and 10 covariates. The experiments are performed using 100 simulations, each using a different random seed and using the $\alpha$ value of $0.10$. Additionally, for our proposed model, we evaluate the probabilistic calibration of the predictive distribution. The CCT and CMC models use normalized split WCPS. The results are shown in Figure  2(a) and 2(b).
Data generating process
This data generation process entails both Setup A and Setup B. For Setup A, where $\gamma=1$, represents an experiment where the treatment has no effect, while for Setup B, where $\gamma=0$, represents a heterogeneous treatment effect. X is sampled from a uniform distribution, $d$ is the number of covariates, $\tau$ represents the treatment effect, $\gamma$ is the treatment heterogeneity, $\pi$ is the propensity, $\epsilon_{0}$ and $\epsilon_{1}$ are Gaussian noise terms, and $T$ is the treatment assignment.
$X\sim uniform([0,1],size=n*d)$ (12)
$\tau=\frac{2}{1+exp(-12(x_{0}-0.5))}\frac{2}{1+exp(-12(x_{1}-0.5))}$ (13)
$\tau_{0}=\gamma\cdot\tau$ (14)
The propensity $\pi$ is sampled from a Beta distribution with shape parameters 2 and 4.
$\pi=\frac{1+Beta(x_{0},2,4)}{4}$ (15)
$\epsilon_{1}\sim\mathcal{N}(0,1)$ (16)
$\epsilon_{0}\sim\mathcal{N}(0,1)$ (17)
$y_{0}=\tau_{0}+\epsilon_{0}$ (18)
$y_{1}=\tau+\epsilon_{1}$ (19)
The treatment is sampled from a uniform distribution between 0 and 1. If the value is below the propensity the assignment is treatment, otherwise control.
$T\sim(uniform(0,1)<\pi)$ (20)
$y=T\cdot y_{1}+(1-T)\cdot y_{0}$ (21)
D.1.2 Nie and Wager Synthetic Data Experiments
The second synthetic data experiments use the setups of Nie and Wager [5]. Four setups are used: A, B, C, and D. Each setup uses 5000 samples and is split into a training and testing set using a 50% split. Every setup is simulated with uncorrelated noise terms, uses five covariates, and a standard deviation of the error term of 1. The same approach is taken here as in Section D.1.1 with 100 simulations, each using different random seeds and using the same $\alpha$ values. The CCT and CMC models also use normalized split WCPS. The results are shown in Figure 8,  9, and  10.
Besides the aforementioned experiments with the data-generating processes proposed by Nie and Wager [5], we also perform a noise dependency impact experiment with it. This experiment will investigate the impact of changing the dependency between the noise variables, $\epsilon_{0}$ and $\epsilon_{1}$, of $y_{0}$ and $y_{1}$, respectively, on the different conformal frameworks. The only difference with the previous experiment is that the noise correlation coefficient $c$ changes from $-1$ to $1$ in increments of $0.1$. The results are shown in Figure 3.
Data generating process: Setup A
Setup A simulates a difficult nuisance component and an easy treatment effect. $p$ is the number of covariates and $n$ the number of samples. X is created as an independent variable, $n$ rows and $p$ columns thus consisting of $[x_{0},x_{1},...,x_{p}]$ each with length $n$. The other variables are the expected outcome $b$, the propensity of receiving treatment $\pi$, $\tau$ the ITE, $w$ the treatment assignment, $y_{0}$ outcome for control, $y_{1}$ for treatment, $\epsilon_{0}$ noise variable for control, $\epsilon_{1}$ noise variable for treatment, and $y$ the true outcome, all with length $n$.
$X\sim uniform([0,1],size=n*p)$ (22)
$b=sin(\pi\cdot x_{0}\cdot x_{1})+2(x_{2}-0.5)^{2}+x_{3}+0.5x_{4}$ (23)
$\pi=max([0.1,0.1,...,0.1],min(sin(\pi\cdot x_{0}\cdot x_{1}),[0.9,0.9,...,0.9])$ (24)
$\tau=\frac{x_{0}+x_{1}}{2}$ (25)
$w\sim Binomial(1,e)$ (26)
$\epsilon_{0}\sim\mathcal{N}(0,\sigma^{2})$ (27)
$\epsilon_{1}=c\cdot\epsilon_{0}+(1-c)\cdot\mathcal{N}(0,\sigma^{2})$ (28)
$y_{0}=b-0.5\tau+\epsilon_{0}$ (29)
$y_{1}=b+0.5\tau+\epsilon_{1}$ (30)
$y=y_{0}(1-w)+y_{1}\cdot w$ (31)
The models only use $y$, $X$, $w$. All the other variables are for evaluation and testing purposes.
Data generating process: Setup B
Setup B mimics a randomized trial. All other variables are calculated similarly as in setup A except for $X,b,\pi,$ and $\tau$.
$X\sim\mathcal{N}(0,1,size=n*p)$ (32)
$b=max(0,x_{0}+x_{1}+x_{2})+max(0,x_{3}+x_{4})$ (33)
$\pi=0.5$ (34)
$\tau=x_{0}+log(1+exp(x_{0}))$ (35)
Data generating process: Setup C
Setup C mimics a situation with confounding with an easy propensity and a difficult baseline. All other variables are calculated similarly as in setup B except for $b,\pi,$ and $\tau$.
$b=2log(1+x_{0}+x_{1}+x_{2})$ (36)
$\pi=\frac{1}{1+exp(x_{0}+x_{1}+x_{2})}$ (37)
$\tau=1$ (38)
Data generating process: Setup D
Setup D mimics a situation with unrelated treatment and control groups. All other variables are calculated similarly as in setup B except for $b,e,$ and $\tau$.
$b=\frac{max(0,x_{0}+x_{1}+x_{2})+max(0,x_{3}+x_{4})}{2}$ (39)
$e=\frac{1}{1+exp(-x_{0})+exp(-x_{1})}$ (40)
$\tau=max(0,x_{0}+x_{1}+x_{2})-max(0,x_{3}+x_{4})$ (41)
D.2 Semi-Synthetic Experiment
The learners are also evaluated on semi-synthetic datasets following the experiments of Alaa et al. [15]. Three semi-synthetic datasets are used: IHDP [10], NLSM [33], and ACIC2016[31] dataset. The simulations run 100 times with $\alpha=0.1$, 100 MC samples, and with split WCPS using the residual as a conformity measure. Since these datasets do not provide propensity scores, we perform logistic regression to estimate it. The results are found in Figure 4 and  11.
D.2.1 IHDP
The Infant Health and Development Program (IHDP) is a randomized controlled trial to evaluate the effect of home visits from physicians on the cognitive test scores of premature infants [10]. This dataset has 25 covariates with 6 continuous and 19 binary. Using this dataset, a semi-synthetic benchmark was created analogous to the study by Hill, using setup B of their work [10]. The used datasets are provided by Shalit et al. [34]. The final dataset consists of 747 samples, 139 are treated samples and 608 of them are control samples. The data generation process of the semi-synthetic data is as follows:
$Y_{0}\sim\mathcal{N}(exp(X+W)\beta,1)$ (42)
$Y_{1}\sim\mathcal{N}(exp(X\beta)-\omega,1)$ (43)
The $\omega$ is always chosen such that ATT (Average Treatment effect on the Treated) is set to 4. The coefficients $\beta$ are chosen at random from
$(0,0.1,0.2,0.3,0.4)$
each with a probability of
$(0.6,0.1,0.1,0.1,0.1)$
respectively. X represents the covariates in the IHDP dataset and W is an offset matrix. The same treatment variable is kept as in the original IHDP study.
D.2.2 NLSM
The National Study of Learning Mindsets (NLSM) is a large-scale randomized trial to investigate the effect of behavioural intervention on the academic performance of students [33]. Using the original NLSM dataset, new data is synthetically generated that follow the same distributions as the NLSM dataset as elaborated in Carvalho et al. [33], Alaa et al. [15]. The resulting dataset contains 10000 data points and 11 covariates ($X=[x_{0},x_{1},...,x_{10}]$). The data generation process for the semi-synthetic data is as follows:
$\gamma\sim\mathcal{N}(0,1,size=76)*0.105$ (44)
$\tau=0.228+0.05\cdot\mathcal{I}(x_{6}<0.07)-0.05\cdot\mathcal{I}(x_{7}<-0.69)-0.08\cdot\mathcal{I}(x_{4}\in[1,13,14]+\gamma_{x_{0}-1}$ (45)
with $\mathcal{I}$ the indicator function and $\gamma_{x_{0}-1}$ the $(x_{0}-1)$th element of an array of Gaussian samples with size 76. Then the outcome is defined as follows:
$Y=W(Y-\tau)+(1-W)*\tau$ (46)
with W as the treatment variable.
D.2.3 ACIC2016
The dataset from the 2016 Atlantic Causal Inference Conference was created for a causal inference competition [31]. It includes 4,802 observations with 58 covariates (3 categorical, 5 binary, 27 counts, and 23 continuous). The competition used these covariates to generate 77 different settings, each with varying outcome surface complexity, degrees of confounding, overlap, and treatment effect heterogeneity. We standardized all covariates and, for each setting, conducted 10 simulation runs, with a test set comprising 20% of the data, randomly generated, following the methodology of Curth and van der Schaar [35], Alaa et al. [15].
Appendix E Experiments Results
Figure 5: Simulation results for setups A and B from Alaa et al. [15] for 90% ITE prediction interval. (With other CMC meta-learners)
Figure 6: Simulation results for setups A, B, C, and D from Nie and Wager [5], with increasing correlation between the noise terms of potential outcome under treatment $Y^{1}$ and control $Y^{0}$. The vertical lines represent the variation in the different simulation runs and cover an interval with the results of 95% of the runs.
Figure 7: Simulation results for setups A, B, C, and D from Nie and Wager [5], with changing correlation between the noise terms of potential outcome under treatment $Y^{1}$ and control $Y^{0}$. The vertical lines represent the variation in the different simulation runs and cover an interval with the results of 95% of the runs.
Figure 8: Simulation results for setups A, B, C, and D from Nie and Wager [5] for estimating 90% ITE prediction interval.
Figure 9: Simulation results for setups A, B, C, and D from Nie and Wager [5] evaluating probabilistic calibration (PIT plot) of CCT- and CMC T-learner with WCPS.
Figure 10: Simulation results for setups A, B, C, and D from Nie and Wager [5] evaluating probabilistic calibration (PIT plot) of CCT- and CMC T-learner with unweighted CPS.
Figure 11: Simulation results for estimating 90% ITE prediction interval on the IHDP and NLSM semi-synthetic datasets.